This is page 2/2. 1 This is page 3/3 with a short special. Also, a rule.

2

This is page 4/4 with a long special. Note the control character; this is meant to confuse dvii when it cuts of the long special text.

This is page 5/5 with 5 specials. 4

This is page 6/-1 with font cmr10 scaled 1200.

i This page has a special with control characters. (Oooh!) Followed by several blank pages.

iii